# Predictable Compression Failures Why Language Models Actually Hallucinate

### v1

@@ 3.5 Operational Planners
$\text{B2T}(x;h^{*})=\operatorname{KL}(\operatorname{Ber}(1-h^{*})\|\operatorname{Ber}(q_{\text{lo}}(x)))$ $\text{RoH}(x)=1-p_{\max}(\bar{\Delta}(x),\bar{q}(x))$ $\text{ISR}(x)=\bar{\Delta}(x)/\operatorname{B2T}(x;h^{*})$ Decision rule: ISR < 1 → abstain; ISR $\geq 1$ → answer
@@
Box 2: Worked example: From EDFL to a decision
Setup. Target reliability $h^{*}=0.05$ (i.e., $p=0.95$). Compute B2T for several conservative priors $q_{\text{lo}}$:
$\mathrm{B2T}(h^{*}=0.05;\,q_{\text{lo}})=\mathrm{KL}(\mathrm{Ber}(0.95)\,\|\,\mathrm{Ber}(q_{\text{lo}})).$ Numerics (nats): $\begin{array}[]{c|ccc}q_{\text{lo}}&0.02&0.10&0.30\\ \hline\cr\mathrm{B2T}&3.519&1.994&0.963\end{array}$ Budget $\bar{\Delta}$ to reliability. Given $\bar{q}=0.10$, the maximum achievable success at budget $\bar{\Delta}$ solves $\mathrm{KL}(\mathrm{Ber}(p)\,\|\,\mathrm{Ber}(0.10))\leq\bar{\Delta}$. Selected budgets $\to$ $p_{\max}$: $\begin{array}[]{c|cccc}\bar{\Delta}\ (\text{nats})&0.5&1.0&2.0&3.0\\ \hline\cr p_{\max}&0.495&0.689&0.951&\approx 1.000\end{array}$ ISR gate. If $\bar{q}=0.10$ and measured $\bar{\Delta}=2.0$, then $\mathrm{ISR}=\bar{\Delta}/\mathrm{B2T}=2.0/1.994\approx 1.00\Rightarrowanswer$. If instead $q_{\text{lo}}=0.02$, then $\mathrm{B2T}=3.519$ and the same budget gives $\mathrm{ISR}\approx 0.57\Rightarrowabstain$ or acquire info.
Algorithm 1 ISR gating with permutation mixture
1:Prompt $x$, target reliability $h^{*}$, permutations $m$, clip $B$, predicate $g$
2:Sample permutations $\{\pi_{k}\}_{k=1}^{m}$ of evidence; form inputs $\Gamma_{\pi_{k}}(x)$
3:for $k=1$ to $m$ do
4:  Query model $\to$ predictive $S_{k}(\cdot)=S(\cdot\mid\Gamma_{\pi_{k}}(x))$
5:  Compute prior term $q_{k}:=S_{k}(g(Y){=}1)$
6:  Compute budget term $u_{k}(y):=\log\frac{P(y)}{S_{k}(y)}$ for reference $P$ (or an estimator)
7:end for
@@
12:if $\mathrm{ISR}\geq 1$ then
13:  return Answer (generate with guardrails)
14:else
15:  return Abstain (or acquire information and re-evaluate)
16:end if
@@ 4 Experiments
Experimental setup. We use our custom-built Factuality Slice dataset comprising 3,059 evidence-grounded QA items from FEVER, HotpotQA, NQ-Open, and PopQA with controlled evidence chunks and hard negatives (see Appendix H for complete documentation). All prompts, scoring code, and fixed seeds used in this paper are documented in Appendix H to permit exact replication without hyperparameter tuning.
4.1 Experiment 1: MDL Optimality via Permutation Mixtures
4.1.1 Ground-Truth Validation
@@
We conduct large-scale dispersion studies on 3,059 binary classification items from all splits spanning $n\in[3,60]$ chunks (58 distinct $n$ values). Each item contains 48-token-capped evidence chunks with binary gold labels. We test Qwen2-7B-Instruct (m=12 unique banded permutations) and Llama-3.1-8B-Instruct (m=16 unique banded permutations), both with 4-bit NF4 quantization. We draw unique banded permutations (6 bands, shuffle within), compute predictions $q_{\pi}(x)=P_{\pi}(\text{"1"})/(P_{\pi}(\text{"1"})+P_{\pi}(\text{"0"}))$ via renormalized label token probabilities, and form uniform mixture $\bar{q}(x)=\frac{1}{m}\sum_{k=1}^{m}q_{\pi_{k}}(x)$.
@@ Results.
Mean absolute residual $|R_{\pi}|$ follows $a+b\ln n$ across both models, confirming Theorem 1. Qwen2-7B shows stronger positional sensitivity than Llama-3.1-8B, reflecting architectural differences. Mean absolute residuals remain approximately 69% of $E_{\text{pair}}$ at $n=60$ for both models.
Model Qwen2-7B Llama-3.1-8B
Dispersion Slope $b$ (vs. $\ln n$) 0.377 [0.319, 0.435] 0.147 [0.109, 0.184]
@@
4.1.2 Self-Consistency Analysis
@@
4.2 Experiment 2: Causal Dose-Response Analysis
To establish causality, we conducted randomized experiments controlling information content while holding prompt length constant at L=4 chunks. We vary the dose $d\in\{0,1,2,3\}$ of support chunks versus non-support chunks.
Answer rate increases by 37.5 percentage points (pp), accuracy by 45.6pp, and hallucination decreases by 17.6pp from dose 0 to 3. $\bar{\Delta}$ increases at 0.375 nats per dose (Spearman $\rho=0.80$, p<0.001). The critical threshold where accuracy exceeds hallucination occurs at dose approximately equal to 1. Using random within-band order as exogenous variation, OLS slope of 0.127 fewer hallucinations per nat establishes information insufficiency as causal mechanism. The Llama-3.1-8B validation yields $\beta\approx 0.110$ hallucination reduction per nat (IV robustness in Appendix D). We use symmetric clipping for stability in the main experiments; min-clipping provides a provable lower-bound (Appendix A.5).
4.3 Experiment 3: Pre-specified audit (deployment calibration)
We evaluate on Gemma-2-9B fine-tuned on FEVER, HotpotQA, NQ-Open, and PopQA with 528 held-out items. Pre-specified settings: evidence permutations with seeds $\{0,1,\ldots,5\}$, symmetric clipping at B=6 nats, preservation of role markers. Information budget: $\bar{\Delta}=\frac{1}{m}\sum_{k=1}^{m}\operatorname{clip}(\log P(y)-\log S_{k}(y),B)$.
Results: 96.2% boundary alignment [94.3, 97.5], near-0% hallucination [0.0, 0.7], 24.1% abstention [20.6, 27.9], 80.5% accuracy on attempts [76.8, 83.8]. Mean Jensen Gap $\hat{\mathfrak{J}}_{\Gamma}$ = 0.82 nats [0.71, 0.93]. Parameter sensitivity ($m\in{3,6,12}$, $B\in{4,6,8}$) shows robust alignment (details in Appendix E). Permutation skeleton ablations show that slopes vary from 0.245 ($K_{\text{bands}}$=3) to 0.291 (uniform permutations), with our 6-band configuration yielding intermediate values.
For rare events with low prior mass: $\bar{q}=0.000$ requires 5.29 nats (ISR=0.16→Refuse); $\bar{q}=0.167$ requires 2.48 nats (ISR=1.06→Answer). Models achieve near-0% hallucination through calibrated refusal rather than random guessing. We find $m=3$ captures most of the Jensen gain with 3× model calls; escalate to $m=6$ only when ISR<1 for critical decisions. This latency-reliability tradeoff enables practical deployment with controlled computational overhead.
@@ 5 Synthesis
The three experiments provide complementary validation, each mapping directly to our theoretical contributions. Experiment 1 validates Theorem 1 (QMV) through observed $O(\log n)$ scaling and Theorem 4 (MDL optimality) through strictly positive Jensen gaps. Experiment 2 validates Theorem 5 (EDFL) by showing causal dose-response of hallucination to information budgets. Experiment 3 demonstrates the operational planners derived from EDFL work in practice with ISR=1.0 cleanly separating safe answering from abstention.
@@ 6 Discussion
6.1 Theoretical Implications
Our framework resolves the fundamental paradox through the key insight that positional encodings induce minimization of $\mathbb{E}_{\pi}[\ell(Y|\Gamma_{\pi}(X))]$ rather than $\ell(Y|X)$, which admits a Kolmogorov-complexity interpretation up to additive constants. The theoretical results are modular and architecture-agnostic:
1. Local stability suffices for logarithmic scaling. The adjacent-swap stability assumption (Assumption 1) is satisfied by a broad class of positional encodings (RoPE, ALiBi, learned) because LayerNorm, bounded Q/K norms, and softmax Lipschitz naturally imply small changes under adjacent swaps. Our bound controls an average-over-permutations quantity in the logit; it is complementary to analyses that consider worst-case positional deviations.
2. Assumption-free certificates provide safety guarantees. The JS-based bound (Proposition 2) requires no structural assumptions and can be computed directly from model outputs, providing per-item safety certificates.
3. Hallucinations are predictable compression failures. Rather than stochastic errors, hallucinations arise deterministically when information budgets fall below EDFL thresholds.
4. Calibrated abstention emerges from information-theoretic principles. Models naturally refuse when ISR < 1, suggesting safe behavior emerges from proper information management.
Our goal is a normative, information-theoretic account with deployable planners. The empirical sections are illustrative checks of the theory’s key predictions (dispersion scaling, mixture gains, and dose-response), not a benchmark study. We therefore do not add detector baselines or coverage-risk curves; these evaluate a different objective (post-hoc detection performance) than the ex-ante budgeting question we target. We release derivations and implementation details so that future work can plug EDFL/ISR into any evaluation suite.
6.2 Practical Applications
Our operational planners enable: pre-generation risk assessment via ISR calculation; graduated responses using ISR thresholds (ISR < 0.5: refuse, 0.5 to 1.0: hedge, >1.0: answer); information acquisition when ISR < 1; and ensemble robustness with $m\geq 6$ permutations. The ground-truth experiments demonstrate permutation averaging provides practical benefits with reasonable computational overhead.
A practical training-time regularizer directly targets the positional Jensen penalty: $\mathcal{L}+\lambda\mathbb{E}_{x}\mathbb{E}_{\pi,\pi^{\prime}}\left[(\operatorname{logit}q_{\pi}(x)-\operatorname{logit}q_{\pi^{\prime}}(x))^{2}\right]$, shrinking prediction variance across permutations without changing the core architecture. This approach complements our inference-time methods.
@@ 6.3 Limitations
We restrict evaluation to binary adjudication because EDFL’s guarantees are tightest for Bernoulli events. While EDFL extends to multi-class via one-vs-rest, the theory is sharpest for Bernoulli predicates that govern abstain/answer decisions in deployment. We therefore focus on Bernoulli adjudication and leave comprehensive multi-class evaluations to follow-on work. For structured outputs (code generation, reasoning chains), compositional predicates like unit tests or rubric satisfaction provide natural binary signals. The relationship between model scale and positional bias (Qwen2 vs Llama variation) deserves systematic investigation across architectures.

### v2

@@ 3.5 Operational Planners
EDFL uses $p$ to denote the (unknown) achieved success probability of an event under the reference distribution. In contrast, planning specifies a target success probability $p^{\star}=1-h^{\star}$, and B2T/ISR are computed relative to this target.
@@
$\text{B2T}(x;p^{\star})=\operatorname{KL}(\operatorname{Ber}(p^{\star})\|\operatorname{Ber}(q_{\text{lo}}(x)))$ $\text{RoH}(x)=1-p_{\max}(\bar{\Delta}(x),\bar{q}(x))$ $\text{ISR}(x)=\bar{\Delta}(x)/\operatorname{B2T}(x;p^{\star})$ Target: set $p^{\star}=1-h^{\star}$. Decision rule: ISR < 1 → abstain; ISR $\geq 1$ → answer
@@
A worked numerical example translating EDFL into B2T/RoH/ISR and an answer/abstain decision appears in Appendix B.3.
Algorithm 1 ISR gating with permutation mixture (Bernoulli instantiation)
0: Prompt $x$, target hallucination rate $h^{\star}$ (so $p^{\star}=1-h^{\star}$), permutations $m$, clip $B$, predicate $g$, reference $p_{\mathrm{ref}}:=P_{\mathrm{ref}}(g(Y){=}1)$
1: Sample permutations $\{\pi_{k}\}_{k=1}^{m}$ of evidence; form inputs $\Gamma_{\pi_{k}}(x)$
2:for $k=1$ to $m$ do
3:  Query model $\to$ $q_{k}:=S(g(Y){=}1\mid\Gamma_{\pi_{k}}(x))$
4:  Compute budget term $u_{k}:=\operatorname{KL}(\operatorname{Ber}(p_{\mathrm{ref}})\parallel\operatorname{Ber}(q_{k}))$
5:end for
@@
10:if $\mathrm{ISR}\geq 1$ then
11:  return Answer (generate with guardrails)
12:else
13:  return Abstain (or acquire information and re-evaluate)
14:end if
In our supervised experiments, $P_{\mathrm{ref}}$ is induced by gold labels and $p_{\mathrm{ref}}$ is a point mass (with $\varepsilon$-smoothing), so $u_{k}$ reduces to the usual negative log-likelihood under each permutation. In deployment, $P_{\mathrm{ref}}$ should be read as a specification oracle / verification primitive for the predicate $g$.
3.5.1 Deployment instantiations (operationalizing $P_{\mathrm{ref}}$)
Algorithm 1 requires a Bernoulli reference $p_{\mathrm{ref}}$ for the predicate $g$. This is not available pre-answer in many open-world grounded QA settings. The intended deployments are those where the predicate is verifiable either (i) before generating long-form content (adjudication or tool-backed decisions), or (ii) after drafting but before release (post-hoc release filtering/retry). Three concrete instantiations are:
  1. (i)
    Binary adjudication with a deterministic oracle (pre-answer). Let the system’s output be a binary adjudication $Y\in\{0,1\}$ (e.g., approve/deny, support/refute) and let a rules engine or database lookup provide $y_{\mathrm{ref}}(x)\in\{0,1\}$ at decision time. Set $P_{\mathrm{ref}}=\delta_{y_{\mathrm{ref}}(x)}$, choose $g(y)=y$, so $p_{\mathrm{ref}}=y_{\mathrm{ref}}(x)$, and compute $q_{k}=S(g{=}1\mid\Gamma_{\pi_{k}}(x))=S(Y{=}1\mid\Gamma_{\pi_{k}}(x))$ from label-token probabilities (as in Experiment 1).
  2. (ii)
    Multiple-choice / discrete-tool answers (pre-answer). Let the answer be one of $K$ discrete options and let a tool/KB return the correct option $y_{\mathrm{ref}}(x)\in\{1,\dots,K\}$. Define $g(y)=\mathbb{1}_{\{y=y_{\mathrm{ref}}(x)\}}$, so $p_{\mathrm{ref}}=1$, and compute $q_{k}=S(g{=}1\mid\Gamma_{\pi_{k}}(x))=\sum_{y:\,g(y)=1}S(Y{=}y\mid\Gamma_{\pi_{k}}(x))$ via normalized option-token probabilities.
  3. (iii)
    Post-hoc verifiable generation (pre-release). Let $y$ be a drafted artifact (e.g., code, a structured query) and let a verifier $V(x,y)\in\{0,1\}$ implement the predicate (unit tests, constraint checkers). Apply Algorithm 1 as a release gate: first draft $y$, then set $p_{\mathrm{ref}}:=V(x,y)$ and compute $q_{k}$ by asking a predictor/verifier model for $\Pr[V(x,y)=1\mid\Gamma_{\pi_{k}}(x)]$ under evidence permutations. If the verifier returns a calibrated probability rather than a hard label, use that value as $p_{\mathrm{ref}}$; guarantees are then conditional on verifier calibration.
@@ 4 Experiments
Experimental setup. We evaluate on a five-benchmark evidence-grounded QA suite: four established real-world benchmarks (FEVER, HotpotQA, NQ-Open, PopQA) plus a Controls benchmark (insufficient-evidence and recency-trap items). We standardize these benchmarks into a common evidence-grounded format that we call Factuality Slice; the combined suite contains 3,059 items with controlled evidence chunks and hard negatives (Appendix H). Experiment 1 studies permutation dispersion and mixture gains under binary labels; Experiments 2–3 evaluate answer/abstain decisions.
Meta-benchmark assembly. We standardize each benchmark into the same evidence-grounded schema (question, evidence chunks with provenance, support spans, and hard negatives), with fixed chunking/capping and BM25-based negative mining. The suite is reproducible by construction: we pin SHA-256 hashes for all raw downloads and use fixed-seed train/val/test splits; Appendix H documents inclusion criteria, licensing, and build details.
Metrics. We report coverage (1–abstention), accuracy on attempts (task-correct among answered items), and hallucination rate (fraction of answered items with $g(Y)=0$) for the binary adjudication predicate $g$ (Appendix B); these need not coincide, so we report both. For the held-out audit we also report boundary alignment: the fraction of items for which the ISR decision agrees with adjudication (answer when g=1g{=}1, abstain when g=0g{=}0). Uncertainty, compute, and reproducibility conventions are summarized in Section 4.1; dataset build and licensing details are in Appendix H.
4.1 Experimental reporting and reproducibility
Uncertainty.
For proportions (coverage/abstention, hallucination rate, accuracy on attempts, boundary alignment) we report 95% Wilson score intervals over items. For the dispersion slopes $b$ in Table 2, we fit OLS to the 58 per-$n$ mean residuals and report the standard 95% confidence interval for the slope coefficient. Unless explicitly shown, other quantities are point estimates.
Compute/resources.
We report algorithmic compute via the total number of model forward passes (one forward pass per prompt/permutation to obtain next-token label probabilities). Evidence chunks are capped at 48 tokens, so the evidence portion ranges from $3\times 48$ to $60\times 48$ tokens (plus question/instruction text).
Table 1: Compute summary (algorithmic). Total forward passes are computed as $\text{items}\times m\times\text{models}$, where $m$ is the number of evaluated evidence permutations per item.
Experiment Models Items $\times\,m$ Total fwd passes
Exp. 1 (disp./mix.) 2: Qwen2-7B Llama-3.1-8B $3{,}059\times 16$ 97,88897{,}888
Exp. 3 (audit) Gemma-2-9B FT $528\times 6$ 3,1683{,}168
Reproduction path.
To preserve double-blind review, URLs for code and processed datasets are withheld in the main paper; we provide full methodological detail in the paper/appendix and will release the dataset builder and experiment scripts at camera-ready. Fine-tuning details for the Gemma-2-9B audit model (Experiment 3): LoRA with r=32r{=}32, $\alpha{=}64$, dropout 0.05 targeting attention and MLP projection modules; BF16 with FlashAttention-2; max sequence length 2048; gradient accumulation 64 with per-device batch size 1; 1 epoch; learning rate $1.5\times 10^{-5}$; weight decay 0.1; cosine schedule with 0.03 warmup ratio; seed 42.
4.2 Experiment 1: Dispersion scaling and mixture gains
Question.
How large is permutation-induced dispersion under exchangeable evidence, and do uniform permutation mixtures yield the expected Jensen gains (and near-optimal mixtures) in practice?
4.2.1 Dispersion scaling (QMV)
@@
We conduct large-scale dispersion studies on 3,059 binary classification items from all splits spanning $n\in[3,60]$ chunks (58 distinct $n$ values). Each item contains 48-token-capped evidence chunks with binary gold labels. We test Qwen2-7B-Instruct and Llama-3.1-8B-Instruct, both with 4-bit NF4 quantization. For each item we draw $m=16$ unique banded permutations (6 bands, shuffle within), compute predictions $q_{\pi}(x)=P_{\pi}(\text{"1"})/(P_{\pi}(\text{"1"})+P_{\pi}(\text{"0"}))$ via renormalized label token probabilities, and form the uniform mixture $\bar{q}(x)=\frac{1}{m}\sum_{k=1}^{m}q_{\pi_{k}}(x)$.
Metric.
We measure mean absolute residual $|R_{\pi}|=\mathbb{E}_{\pi}|q_{\pi}(x)-\bar{q}(x)|$ and fit the empirical law $|R_{\pi}|\approx a+b\log n$ (Figure 1, Table 2).
@@ Results.
Mean absolute residual $|R_{\pi}|$ follows $a+b\log n$ across both models, confirming Theorem 1. Qwen2-7B shows stronger positional sensitivity than Llama-3.1-8B, reflecting architectural differences. Mean absolute residuals remain approximately 69% of $E_{\text{pair}}$ at $n=60$ for both models.
Table 2: Summary of permutation dispersion and mixture effects (Experiment 1). N=3,059N=3{,}059 items; banded permutations with $m=16$ per item and $n\in[3,60]$ chunks (48-token cap). Bracketed intervals for $b$ are 95% CIs as defined in Section 4.1.
Metric Qwen2-7B Llama-3.1-8B
Dispersion Slope $b$ (vs. $\log n$) 0.377 (0.319, 0.435) 0.147 (0.109, 0.184)
@@
4.2.2 Mixture gains and near-optimality
@@
4.3 Experiment 2: Causal dose-response of information
Question.
Does increasing information content causally reduce hallucination when prompt length is held fixed?
Design.
To identify the effect of information content (not prompt length), we run randomized experiments holding the prompt length fixed at $L=4$ chunks while varying the dose $d\in\{0,1,2,3\}$ of support chunks versus non-support chunks.
Identification.
Dose randomization shifts the information budget $\bar{\Delta}$ while holding length constant. Appendix D reports robustness checks (including IV using $d$ as an instrument for $\bar{\Delta}$).
Results.
From dose 0 to 3, answer rate increases by 37.5 percentage points (pp), accuracy by 45.6pp, and hallucination decreases by 17.6pp. The information-budget estimate decreases by 0.375 nats per additional support chunk (Spearman $\rho=-0.80$, p<0.001), and the OLS slope is 0.127 fewer hallucinations per additional nat of decrease. A separate Llama-3.1-8B validation yields $\beta\approx 0.110$ hallucination reduction per nat. Figure 2 summarizes these dose-response relationships. We use symmetric clipping for stability; min-clipping provides a provable lower bound (Appendix A.5).
Interpretation.
Because dose shifts $\bar{\Delta}$ while holding length constant, the negative dose-response supports the EDFL mechanism: information sufficiency improves with supportive evidence and reduces hallucination risk under Bernoulli adjudication (Theorem 5).
4.4 Experiment 3: Pre-specified held-out audit
Question.
Does the fixed ISR=1=1 boundary generalize out-of-sample as a reliability–coverage operating point without threshold tuning?
Design.
We evaluate a Gemma-2-9B model fine-tuned on the four non-control benchmarks in Factuality Slice (FEVER, HotpotQA, NQ-Open, PopQA) on a 528-item held-out audit drawn from the full five-benchmark suite (including Controls). We pre-specify the permutation seeds $\{0,1,\ldots,5\}$ (so $m=6$), preserve role markers, and use symmetric clipping with $B=6$ nats. Boundary alignment is the fraction of items for which the ISR gate’s decision agrees with adjudication (answer when g=1g{=}1, abstain when g=0g{=}0). In supervised evaluation, $P_{\mathrm{ref}}$ is induced by the gold label, so the budget reduces to $\bar{\Delta}=\frac{1}{m}\sum_{k=1}^{m}\operatorname{clip}(-\log q_{k},B)$ where $q_{k}=S_{k}(g{=}1)$ is the model probability assigned to the correct outcome.
Results.
Boundary alignment is 96.2% [94.3, 97.5], hallucination rate is 0.0–0.7%, abstention is 24.1% [20.6, 27.9], and accuracy on attempts is 80.5% [76.8, 83.8]. The mean Jensen gap is $\hat{\mathfrak{J}}_{\Gamma}=0.82$ nats. Figure 3 reports the corresponding reliability–coverage operating point.
Interpretation and robustness.
[…]
@@ 5 Synthesis
The three experiments provide complementary validation, each mapping directly to our theoretical contributions. Experiment 1 validates Theorem 1 (QMV) through observed $O(\log n)$ scaling and Theorem 4 (MDL optimality) through strictly positive Jensen gaps. Experiment 2 validates Theorem 5 (EDFL) by showing causal dose-response of hallucination to information budgets. Experiment 3 demonstrates the operational planners derived from EDFL work in practice with an ISR=1=1 operating point separating safe answering from abstention.
@@ 6 Discussion
Implications.
In grounded QA, evidence order acts as a nuisance variable: different permutations induce measurable dispersion in predictions (Experiment 1; Figure 1). QMV explains why dispersion can grow with context depth under mild positional sensitivity (Theorem 1 and Theorem 2), while permutation mixtures recover log-loss improvements consistent with Jensen’s inequality. EDFL links reliability to information budgets for Bernoulli adjudication (Theorem 5); randomized dose-response supports that increasing information reduces hallucination (Experiment 2; Figure 2).
Operational takeaway.
ISR gating yields a concrete abstention interface on the reliability–coverage plane (Experiment 3; Figure 3): when ISR is low, abstain or acquire information; when ISR is high, attempt an answer.
@@ 6.1 Limitations
  • Binary adjudication focus. EDFL is sharpest for Bernoulli predicates; extensions to multi-class or structured generation require careful predicate design.
  • Reference for the predicate. In evaluation, $P_{\mathrm{ref}}$ comes from gold labels; in deployment, it comes from the system’s verification primitive or a calibrated verifier. Guarantees are relative to $P_{\mathrm{ref}}$ and must account for verifier error.
  • Compute overhead. Permutation mixtures require multiple model calls ($m$), trading latency for reliability.
  • Permutation scheme dependence. Banded vs uniform permutations (and chunking choices) can change dispersion constants; we partially probe sensitivity in Appendix E.
  • Scope of evidence grounding. Results are established on an evidence-grounded QA dataset; generalization to open-ended generation is not claimed.
6.2 Broader impact and ethics
Potential positive impacts. ISR-style gating can make abstention a first-class, auditable decision in evidence-grounded QA. Potential negative impacts. The same techniques increase compute (more model calls) and could be misused outside their assumptions; we emphasize scope limits (Section 6.1) and treat the verifier as part of the system that must be audited.